OpenStudy (anonymous):

Which medium (the glass, water or diamond), is the most dense and how do you know?

OpenStudy (shayaan_mustafa):

Hi jacquelyne :) How are you dear? Welcome to Openstudy :D

OpenStudy (shayaan_mustafa):

The more the refractive index (\( \eta\)) the more the medium density. \(\eta\text{ of water }\)=1.33 \(\eta\text{ of glass }\)=1.6 approx \(\eta\text{ of diamond }\)=2.419 approx 2.42 So ofcourse then Diamond is the most dense.

OpenStudy (anonymous):

There is a more direct way of determining that. Density is defined as mass/volume, so all you have to do is take a known volume of those materials and weigh them.

OpenStudy (shane_b):

I wouldn't go by the refractive index because it's based on the propagation of light....and materials which are not dense can still have a high refractive index. Use the density formula:\[density=\frac{mass}{volume}\] Common density values for diamond, water and glass are: \[d_{diamond}=\frac{3250kg}{m^3}\]\[d_{water}=\frac{1000kg}{m^3}\]\[d_{glass}=\frac{2600kg}{m^3}\]